OpenStudy (anonymous):

Add and simplify if possible: x+3/x-4 + x/4-x

jimthompson5910 (jim_thompson5910):

Hint: 4-x = -1(x-4) So \[\Large \frac{x+3}{x-4}+\frac{x}{4-x}\] \[\Large \frac{x+3}{x-4}+\frac{x}{-1(x-4)}\] \[\Large \frac{x+3}{x-4}-\frac{x}{x-4}\]

OpenStudy (anonymous):

\[\frac{2x + 3}{ x-4 }\]

OpenStudy (anonymous):

is that right?

jimthompson5910 (jim_thompson5910):

No, \[\Large \frac{x+3}{x-4}+\frac{x}{4-x}\] \[\Large \frac{x+3}{x-4}+\frac{x}{-1(x-4)}\] \[\Large \frac{x+3}{x-4}-\frac{x}{x-4}\] \[\Large \frac{x+3-x}{x-4}\]

OpenStudy (anonymous):

oh so 3/ x-4

jimthompson5910 (jim_thompson5910):

yes, the final answer is \[\Large \frac{3}{x-4}\]

OpenStudy (anonymous):

yea! thanks :)

jimthompson5910 (jim_thompson5910):

np